# Quintessential interpretation of the evolving dark energy in light of DESI

### v2

@@ II Reconstruction of the scalar potential
The above linear relation (1) should be viewed as a toy model, or the simplest nontrivial parameterization of $w(a)$ with time dependence [63, 64, 65]. The DESI data prefer $w_{a}<0$. Obviously, $w$ can become smaller than $-1$ at an early time and violate the null energy condition. [^2] It was suggested that such a phantom phase is a mere consequence of an inappropriate choice of priors [66], after the appearance of the first version of our paper. A (homogeneous) canonical scalar field $\phi$ with positive potential $V(\phi)\geq 0$ can realize only $-1\leq w\leq 1$, so the interpretation in terms of $\phi$ must break down at some point. On the other hand, $w$ also exits this range in the future extrapolation. This ($w>1$) is associated with a less exotic realization by a negative potential $V(\phi)<0$ in the relevant field domain. We will come back to these points below.
@@
Figure 1: The $1\sigma$ and $2\sigma$ contours of the allowed values of $V$ and $V^{\prime}$ at the present time. The blue, green, and orange contours correspond to Pantheon$+$, DES, and Union, respectively, combined with CMB and DESI.
Let us translate the dynamics of dark energy into the quintessential field $\phi=\phi(t)$. Its EoS parameter is given by $w=\frac{\frac{1}{2}\dot{\phi}^{2}-V}{\frac{1}{2}\dot{\phi}^{2}+V}$. Using the Friedmann equations, the kinetic energy, the scalar potential, and its derivative are given in terms of $w(a(t))$, and $a(t)$ as follows:
$\begin{gathered}\frac{1}{2}\dot{\phi}^{2}=\frac{1}{2}\left(1+w\right)\rho_{\text{DE}},\quad V=\frac{1}{2}\left(1-w\right)\rho_{\text{DE}},\\ V^{\prime}=\frac{1}{2}\left(w_{a}a-3(1-w^{2})\right)H\sqrt{\frac{\rho_{\text{DE}}}{1+w}}.\end{gathered}$ (4)
This can be used to map the contour on the $(w_{0},w_{a})$-plane to the contour on the $(V,V^{\prime})$-plane. Fig. 1 shows the contour evaluated at the present time.
For an intuitive understanding, we show the reconstructed scalar potential $V(\phi)$ in Fig. 2 and the time evolution of $\phi(t)$ as well as $a(t)$ in Fig. 3 with the central value of DESI$+$CMB$+$DES ($w_{0}=-0.727$ and $w_{a}=-1.05$) as the benchmark parameter. For this purpose, we define the origin of $\phi$ to coincide with the current value, i.e., $\phi(t_{0})=0$ and assume $\dot{\phi}(t_{0})>0$ without loss of generality.
@@
It is also intriguing to discuss the implications for the future of the Universe. Fig. 3 shows that the accelerated expansion [67, 68] will soon stop and it will turn to the decelerated expansion again. By construction with the linearly increasing $w$, $V$ asymptotes to 00 from below with slowly rolling-up $\phi$. Of course, we can easily imagine that the linear behavior $w(a)$ changes at some point in the future, and the shape of the potential may be modified. If there is a minimum with $V>0$, there will be another accelerated expansion phase in the future with the reduced dark energy. On the other hand, if the field is trapped in a minimum with $V<0$ or if the potential is unbounded below, the Universe will eventually turn around into a contracting phase [69, 70]. In such a case, the kinetic energy of $\phi$ typically dominates the energy density of the Universe and it will lead to a big crunch.

### v3

@@ II Reconstruction of the scalar potential
The purpose of this paper is to interpret the DESI result in terms of quintessence. In this section, we assume exactly the form in Eq. (1) and reconstruct the scalar field dynamics. We consider a (homogeneous) canonical scalar field $\phi$ with its potential $V(\phi)$. In general, the EoS of such a field is given by $w=\frac{\frac{1}{2}\dot{\phi}^{2}-V}{\frac{1}{2}\dot{\phi}^{2}+V}$. For an arbitrary non-negative potential $V(\phi)\geq 0$, the EoS parameter is restricted as $-1\leq w\leq 1$. As is well known, positive $V$ with small kinetic energy realizes $w\simeq-1$, surving as dark energy. Negative potential $V<0$ allows values of $w$ outside of the above range, but a smooth transition from $w\gtrsim-1$ with $V>0$ to $w\lesssim-1$ with $V<0$ is impossible since $w=1$ at $V=0$ unless the kinetic energy vanishes simultaneously. Sometimes, one considers a phantom scalar field, which has the wrong-sign kinetic term, to realize $w<-1$ with $V>0$, but it is either nonunitary or unstable. Even if the phantom dark energy does not couple directly to the Standard-Model particles, they interact with gravity and the theory is not viable [63, 64].
[…]
@@
Figure 1: The $1\sigma$ and $2\sigma$ contours of the allowed values of $V$ and $V^{\prime}$ at the present time. The blue, green, and orange contours correspond to Pantheon$+$, DES, and Union, respectively, combined with CMB and DESI. We set $\Omega_{\text{m,}0}=0.3$ for simplicity.
Let us translate the dynamics of dark energy into the quintessential field $\phi=\phi(t)$. That is, we reconstruct $V(\phi)$ and the associated solution $\phi=\phi(t)$ that reproduces the specific dynamics (1). Using the Friedmann equation,
$3H^{2}M_{\text{Pl}}^{2}=\Omega_{\text{m},0}a^{-3}+(1-\Omega_{\text{m},0})\frac{\rho_{\text{DE}}(a)}{\rho_{\text{DE},0}},$ (4)
the kinetic energy, the scalar potential, and its derivative are given in terms of $w(a(t))$, and $a(t)$ as follows:
$\begin{gathered}\frac{1}{2}\dot{\phi}^{2}=\frac{1}{2}\left(1+w\right)\rho_{\text{DE}},\quad V=\frac{1}{2}\left(1-w\right)\rho_{\text{DE}},\\ V^{\prime}=\frac{1}{2}\left(w_{a}a-3(1-w^{2})\right)H\sqrt{\frac{\rho_{\text{DE}}}{1+w}},\end{gathered}$ (5)
where $V^{\prime}\equiv\mathrm{d}V(\phi)/\mathrm{d}\phi$ is the derivative of the scalar potential. This can be used to map the contour on the $(w_{0},w_{a})$-plane to the contour on the $(V,V^{\prime})$-plane. To this end, we fix the present matter abundance $\Omega_{\text{m},0}=0.3$ and deal with the combinations $V/(3H_{0}^{2}M_{\text{Pl}}^{2})$ and $V^{\prime}/(3H_{0}^{2}M_{\text{Pl}}^{2})$ so that it is not sensitive to the overall scale $H_{0}$. [^4] In the following, we use the same value of $\Omega_{\text{m},0}$ as a representative value unless otherwise specified since the results do not crucially depend on its precise value. Because of the assumed flatness of space, the dark energy density is obtained as $\Omega_{\text{DE}}=1-\Omega_{\text{m}}$. Fig. 1 shows the contour evaluated at the present time.
From Eqs. (5), we obtain $\dot{\phi}(t)$ and $V(t)$. Integrating the former, we obtain $\phi(t)$, whose integration constant is set such that the origin of $\phi$ coincides with the current value, i.e., $\phi(t_{0})=0$. We also assume $\dot{\phi}(t_{0})>0$ without loss of generality. Typically, we find $\phi$ does not turn around, so $\phi(t)$ can be inverted to $t(\phi)$. Thus, one can reconstruct $V(\phi)=V(t(\phi))$. In addition, we obtain $a(t)$ from the Friedmann equation (4). For an intuitive understanding, we show the reconstructed scalar potential $V(\phi)$ in Fig. 2 and the time evolution of $\phi(t)$ as well as $a(t)$ in Fig. 3 with the central value of DESI$+$CMB$+$DES ($w_{0}=-0.727$ and $w_{a}=-1.05$) as the benchmark parameter.
@@
It is also intriguing to discuss the implications for the future of the Universe by extrapolating Eq. (1). Fig. 3 shows that the accelerated expansion [70, 71] will soon stop and it will turn to the decelerated expansion again. Literally assuming Eq. (1) eventually leads to $w(a)\geq 1$. From the EoS of $\phi$, $w=\frac{\frac{1}{2}\dot{\phi}^{2}-V}{\frac{1}{2}\dot{\phi}^{2}+V}$, we see that $V$ must get negative. Further growth of $w$ corresponds to $V$ asymptoting to 00 from below with slowly rolling-up $\phi$ (see the inset of Fig. 2 and Fig. 3). Of course, we can easily imagine that the linear behavior $w(a)$ changes at some point in the future, and the shape of the potential may be modified. If the final or asymptotic value of $V(\phi)$ is positive, there will be another accelerated expansion phase in the future with the reduced dark energy. On the other hand, if the field is trapped in a minimum with $V<0$ or if the potential is unbounded below, the Universe will eventually turn around into a contracting phase [72, 73]. In such a case, the kinetic energy of $\phi$ typically dominates the energy density of the Universe and it will lead to a big crunch. We emphasize again that any statement about $\phi>0$ relies on the extrapolation of Eq. (1).